An oxide is a compound that contains at least one oxygen atom and only one other element in its chemical formula. Using your knowledge of chemistry comment on the chemistry of oxides.

1
Expert's answer
2022-04-19T03:11:56-0400

Basic oxides are oxides that show basic properties in opposition to acidic oxides and that react with water to form a base for example CaO + H2O = Ca(OH)2

react with an acid to form a salt and water which are called neutralization reactions

CuO + H2SO4 = CuSO4 + H2O

Basic oxides can react with acidic oxides

MgO + CO2 = MgCO3


An acidic oxide is an oxide that produces an acidic solution upon addition to water

SO3 + H2O = H2SO4

react with an alkalis to form a salt and water

SO2 + 2NaOH = Na2SO3 + H2O

SiO2 does not react with water


Amphoteric oxides are the those oxides which behave both as acidic oxides as well as basic oxides. e.g aluminium oxide and zinc oxide.

Amphoteric oxides dissolve in water to form alkaline solutions. Alkaline solutions contain hydroxide ions. Thus aluminium oxide (Al2O3) reacts with hydrochloric acid to form aluminium chloride and water. With sodium hydroxide solution it forms sodium aluminate (NaAlO2) and water.  

ZnO + 2HCl = ZnCl2 + H2O

Al2O3 + 2NaOH = 2NaAlO2 + H2O